OpenStudy (anonymous):

lim x->1 (x^n-1)/(x^m-1) =?

OpenStudy (anonymous):

You can apply L'Hopital's Rule, since it's of the \(\frac{0}{0}\) form: \[\lim_{x \rightarrow 1}{nx^{n-1} \over mx^m-1}={n \over m}\]

myininaya (myininaya):

\[\lim_{x \rightarrow 1}{nx^{n-1} \over mx^{m-1}}={n \over m} \]

OpenStudy (zarkon):

\[\lim_{x \rightarrow 1}\frac{x^{n}-1}{x^{m}-1}=\lim_{x \rightarrow 1}\frac{(x-1)(x^{n-1}+x^{n-2}+\cdots+x+1)}{(x-1)(x^{m-1}+x^{m-2}+\cdots+x+1)}\] \[\lim_{x \rightarrow 1}\frac{x^{n-1}+x^{n-2}+\cdots+x+1}{x^{m-1}+x^{m-2}+\cdots+x+1}\] \[=\begin{array}{cc}n \text{ terms}\\ \frac{\overbrace{1+1+\cdots+1}}{\underbrace{1+1+\cdots+1}}\\m \text{ terms}\end{array}={n \over m}\]

OpenStudy (anonymous):

Nice Zarkon!